OpenStudy (anonymous):

PLEASE HELP, WILL GIVE A MEDAL. Given: 2x + 3y = 6. What is the y-intercept? A. (0, 3) B. (2, 0) C. (0, 2) D. no intercept

TheSmartOne (thesmartone):

Change it to y=mx+b where b is they y-intercept

OpenStudy (anonymous):

hmm, wouldn't it be C, (0, 2) because 2*0 is 0 + 3*2 is 6

TheSmartOne (thesmartone):

So 3y=-2x+6 y=-2/3x+6/3 y=-2/3x+2 So yes it would be (0,2) @vanesilva496

TheSmartOne (thesmartone):

Oops I though you asked the question. @vanesilva496 @vanesilva496 Please refrain from giving direct answers.

TheSmartOne (thesmartone):

@z053sme

OpenStudy (anonymous):

Srry, this is actually my first time helping. It wont happen next time srry

TheSmartOne (thesmartone):

Its okay. You can read this tutorial for new OS users that I made. xD http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

TheSmartOne (thesmartone):

Also read the CoC so you won't get warned or suspended. http://openstudy.com/code-of-conduct